Determine the change in volume of a block of a cast iron 5cm x 10cm x 6cm, when the temperature changes from 15°C to 47°C. The coefficient of linear expansion of the cast iron is
0.000 010°C^-1.
1
Expert's answer
2020-11-12T09:33:45-0500

We can find the change in volume of a block of a cast iron from the formula:


"\dfrac{\Delta V}{V}=\alpha_V\Delta T,"

here, "V=300\ cm^3" is the volume of the block of the cast iron, "\Delta V" is the change in volume of the block of the cast iron, "\alpha_V=3\alpha_L" is the volume thermal expansion coefficient of the cast iron, "\alpha_L=1.0\cdot10^{-5}\ ^{\circ}C^{-1}" is the linear thermal expansion coefficient of the cast iron and "\Delta T=47\ ^{\circ}C-15\ ^{\circ}C=32\ ^{\circ}C" is the change in temperature.

Then, from this formula we can calculate "\Delta V":


"\Delta V=3\alpha_LV\Delta T,""\Delta V=3\cdot 1.0\cdot10^{-5}\ ^{\circ}C^{-1}\cdot 300\ cm^3\cdot 32\ ^{\circ}C=0.29\ cm^3."

Answer:

"\Delta V=0.29\ cm^3."